OpenStudy (anonymous):

A bridge charges 2-axled vehicles a $5 toll and 3-axled vehicles an $8 toll. In an hour the bridge collected $741 from 120 vehicles. If tolls were $1 higher for 2-axled and $2 higher for 3-axled vehicles, how much would the bridge have collected? Show work please. answer choices a) $888 b) $908 c) $926 d) $934 e) $1012

hero (hero):

So x + y = 120 5x + 8y = 741 Solve the system. Once you solve the system and find (x,y) value Calculating this will be easy: 6x + 10y = ?

OpenStudy (anonymous):

Thank you. I had figured it out :)